DCPAM Reference Manual

GFD Dennou Club

平成 17 年 2 月 18 日

目次

1 Module dycore grid mod 2 1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 Future Plans . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.6 Dependency . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.7 Public Interface . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.8 Procedure Interface . . . . . . . . . . . . . . . . . . . . . . . 2

1.8.1 Initialize module . . . . . . . . . . . . . . . . . . . . . 2 1.8.2 Terminate module . . . . . . . . . . . . . . . . . . . . 3

1 Module dycore grid mod

• Developers: Morikawa Yasuhiro

• Version: $Id: dycore grid.f90,v 1.4 2005/01/03 12:48:05 morikawa Exp $

• Tag Name: $Name: $

• Change History:

1.1 Overview

Provide Grid Information.

1.2 Error Handling 1.3 Known Bugs 1.4 Note 1.5 Future Plans 1.6 Dependency

use dycore_type_mod    , only : STRING
use grid_3d_mod        , only : im, jm, km
use grid_wavenumber_mod, only : nm

1.7 Public Interface

private
public :: dycore_grid_init, dycore_grid_end  ! subroutines
public :: im, jm, km, nm                     ! variables

1.8 Procedure Interface

1.8.1 Initialize module

grid 3d mod と grid wavenumber mod の初期化を行なう。

subroutine dycore_grid_init

Dependency

use dycore_type_mod    , only: STRING
use grid_3d_mod        , only: grid_3d_init
use grid_wavenumber_mod, only: grid_wavenumber_init
use dc_trace           , only: DbgMessage, BeginSub, EndSub

1.8.2 Terminate module

grid 3d mod と grid wavenumber mod の終了処理を行なう。

subroutine dycore_grid_end

Dependency

use grid_3d_mod        , only: grid_3d_end
use grid_wavenumber_mod, only: grid_wavenumber_end
use dc_trace           , only: DbgMessage, BeginSub, EndSub